# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 3.3.2 What metrics are used in the evaluation process?
11.11% of the works evaluate their proposals using a single metric [58, 53, 11, 19, 62, 1, 32]. 55.56% of the works use both predictive and ranking quality metrics [28, 29, 43, 17, 15, 21, 81, 67, 70, 77, 66, 80, 30, 54, 56, 49, 50, 4, 25, 57, 74, 20, 75, 45, 18, 60, 26, 65, 23, 64, 3, 55, 68, 69, 63]. 31.75% of the works use only predictive metrics [33, 44, 78, 46, 72, 76, 38, 39, 24, 22, 61, 13, 58, 11, 51, 1, 32, 71, 40, 19]. Lastly, 11.11% of the works employ only ranking metrics [31, 79, 42, 53, 73, 16, 62].
@@
[28, 29]
[63]
[78]
[76, 46, 72, 33, 38, 44]
[43]
[17]
[15, 81, 67, 21]
[80]
[77, 66]
[54, 30, 56]
[13, 39, 22, 61, 24]
[4]
[57, 75, 20, 74, 25, 18]
[23, 26, 64, 65]
[3]
[60, 45]
[32, 1]
[55, 51]
[68, 69]
[31, 79]
[16, 73]
[53]
[11, 19]
[40, 71]
[70]
[42]
[62]
[50]
[49]
[58]
@@
In addition to the previously mentioned metrics, other metrics have also been observed in 25.40% of the works, with Area Under the Curve (AUC) and Hit Rate (HR) being the most commonly used. AUC evaluates the ability of the system to distinguish between relevant and irrelevant papers, while HR is the ratio of the number of relevant documents in the Top-N recommendation list to the total number of relevant documents in the dataset. Another, less known metric, bpref, has been used in one work [63]; it measures retrieval effectiveness by counting how often non-relevant documents are ranked before relevant ones using binary relevance judgments.
@@
AUC [51, 68, 69, 78, 40, 19] 9.52%
HR [55, 16, 70, 73] 6.35%
ACC [78, 71] 3.17%
Success [28, 29] 3.17%
Scalability [80] 1.59%
Recency [80] 1.59%
High citation popularity [80] 1.59%
Diversity [80] 1.59%
True Negative Rate (TNR) [78] 1.59%
MSE [78] 1.59%
bpref [63] 1.59%
@@ 3.3.3 How are the evaluations conducted?
76.19% of the analyzed papers focus on quantitative assessments. 31.25% of them are complemented by both, ablation studies and configuration tests [21, 81, 67, 77, 66, 30, 54, 4, 25, 26, 65, 23, 68, 79, 73]. 29.17% are complemented by ablation studies [28, 29, 70, 80, 56, 45, 18, 64, 55, 78, 24, 71, 40, 16] and 10.42% by configuration tests [49, 20, 39, 51, 57]. Additionally, 8.33% of the works that conduct exclusively quantitative evaluations also assess execution time [25, 71, 57, 33].
14.29% of the works carry out both quantitative and qualitative evaluation, of which 44.44% [43, 69, 75, 63] complement it with some type of ablation study and 44.44% [17, 15, 74, 42] with an ablation study and configuration test.
9.52% of the works conduct only a qualitative evaluation [52, 50, 44, 11, 31, 53], only one includes an additional analysis [50] to evaluate execution time and only 2 conduct online evaluations [15, 44].
@@
[17, 74, 42]
[43, 69, 63]
[60]
[54, 81, 68, 26, 77, 66, 30, 4, 23, 67, 21, 65, 73, 79]
[25]
[16, 28, 29, 80, 64, 55, 78, 40, 70, 56, 18, 24, 45]
[39, 51, 20, 49]
[33]
[76, 22, 3, 13, 58, 61, 32, 1, 46, 72, 19, 38, 62]
[31]
[75]
[15]
[52]
[11]
[57]
[53]
[44]
[50]
[71]
@@
Only 12.70% of the works do not compare their systems or algorithms with others [39, 19, 76, 33, 52, 53, 11, 50], 50% of which rely solely on qualitative evaluations.
@@
Regarding the types of baselines employed, 28.57% of the works utilized a combination of RP-RS and RSs baselines [18, 73, 16, 78, 30, 67, 21, 25, 28, 79, 15, 29, 74, 77, 24, 61, 75, 20], followed by 25.40% that employed only RP-RS baselines [43, 80, 17, 26, 32, 60, 3, 51, 4, 22, 65, 56, 13, 64, 31, 46], and 17.46% that used RSs baselines [1, 40, 68, 81, 66, 45, 69, 55, 70, 72, 57]. Notably, 15.87% of the works did not compare recommendation algorithms but rather focused on paper and user representation algorithms [49, 54, 62, 71, 38, 23, 58, 42, 44, 63]. Although the analyzed time period is relatively brief, the data suggest a potential shift in trend towards the latter.
Regarding datasets, the majority of works (79.37%) use between one and two datasets. More specifically, 47.62% use only one dataset for the evaluation [39, 19, 33, 52, 50, 49, 57, 22, 58, 62, 3, 51, 13, 46, 72, 31, 54, 81, 28, 40, 56, 61, 38, 1, 29, 45, 30, 66, 70, 21], while 31.75% use two datasets [76, 60, 32, 20, 17, 68, 26, 71, 73, 65, 4, 64, 24, 55, 25, 43, 42, 67, 15]. An additional 11.11% use three datasets [16, 18, 77, 80, 79, 69, 63]. It is worth noting that works utilizing more than four or eight datasets to evaluate their results typically rely on a base dataset, building different versions of these to assess specific aspects. In [23], up to four versions of each dataset are generated by randomly selecting 25%, 50%, 75%, and 100% of the data, while in [75, 78], two versions of each base dataset are created to obtain dense and sparse versions.
@@ 3.4 What are the main challenges addressed by the RP-RSs?
Previous surveys in the field identified several open issues in their reviews [9, 6, 35]. In the rapidly evolving field of RP-RSs   significant advances have been made in recent years. However, despite these strides, several challenges persist that hinder the full realization of potential benefits.
@@
The challenge of disregarding user modeling in RP-RSs persists, as many systems rely on generic texts or specific papers provided by users, rather than incorporating personalized user profiles. This approach assumes that the information needs of the user are implicit and not tailored to individual preferences, which limits the overall effectiveness and relevance of recommendations. Personalization, which adapts the recommendations to the unique needs and behaviors of each user, plays a critical role in improving user satisfaction. Compared to [35], the current work shows a 10% increase in the adoption of user modeling, with 61.90% of the reviewed works now addressing this issue and aiming to offer more personalized recommendations.
@@
Relying solely on precision as the primary evaluation metric in RP-RSs overlooks key factors such as recall, ranking, novelty, diversity, and serendipity. While precision measures the accuracy of the top recommendations, recall focuses on completeness, and ranking metrics like NDCG, MAP, and MRR assess how well the system orders relevant papers. However, other important aspects such as novelty (the ability to suggest new and unexpected papers), diversity (offering a broad range of topics), and serendipity (providing surprisingly relevant recommendations) are often under-explored due to their complexity and rarity in use. These factors are crucial for ensuring that recommendations are not only relevant but also engaging and broad in scope, contributing to a richer user experience. Compared to [35], there has been a 25.65% increase in the use of recall and a 18.12% increase in NDCG, but the adoption of less common metrics has decreased by 21.21%, with AUC and HR being the most widely used metrics at 9.52% and 6.35%, respectively.
@@
In the field of RP-RSs serendipity refers to the process of suggesting unexpected yet valuable papers, thereby facilitating the discovery of novel ideas. However, challenges such as balancing novelty with relevance, assessing the impact of serendipitous recommendations, and personalizing the recommendation process have led to serendipity being underemphasized in recent literature, which has predominantly focused on improving accuracy and personalization. Despite these challenges, serendipity holds significant potential in fostering innovation and expanding the scope of research. Among the works reviewed, only one paper explicitly addresses the concept of serendipity [23]. To address this challenge, in [23] the authors propose a citation recommendation framework that integrates multi-topic communities into a modified version of PageRank.
@@
[35] discusses the balance between relevance and diversity in RP-RSs, highlighting that focusing only on relevance can lead to similar recommendations. Diversity in the generated recommendations can better satisfy the needs of the users. The use of MRR as a quality metric reflected the recognition of this issue, as MRR helps ensure that the suggestions are both relevant and diverse. A small number of recent works have specifically targeted the diversity issue in RP-RSs. [49] identifies a new feature capturing the diversity of research interests, moving beyond the recency of publications to model varied researcher interests. [57] introduces a graph of FoS, which uncovers latent knowledge paths to enhance recommendation diversity by discovering relationships in scientific areas. [80] proposes measuring diversity by calculating the sum of shortest paths between pairs of items in the recommendation list, helping quantify diversity. [21] focuses on enhancing diversity by utilizing indirect social ties within academic teams, narrowing down candidate papers based on shared preferences. [15] promotes diversity through a Knowledge Discovery Model (KDM), which evaluates paper relevance by the degree of keyword association, aiming to provide diverse topics and sub-topics. Lastly, [17] employs a heterogeneous network and RippleNet to explore potential connections and expand user interests, enhancing the diversity of the recommendations. In the current survey, a small increment in the use of the MRR metric has been detected, making it the fourth most used metric, with 30.16% of the works employing it.
@@
Many works in RP-RSs fail to provide sufficient documentation, such as details about experimental setups, algorithms, or data sources. This lack of transparency makes it difficult for other researchers to replicate or fully understand the methods and findings. A clear indication that this issue remains unresolved is that 21 works were excluded from further analysis in this survey due to missing essential information. It is worth noting that while [35] did not identify any works sharing their code, this survey has found 6 works that have made their developed code accessible [15, 33, 53, 49, 57, 63].

### v2

@@ 3.3.2 What metrics are used in the evaluation process?
11.11% of the works evaluate their proposals using a single metric [IJ:IEEE:STERLING:2022, IJ:IEEE:SARWAR:2021, IJ:ELECTRONICS:Blazevic:2023, IJ:SHTI:GUO:2022, EC:ECIR:Takahashi:2022, ARXIV:Mohamed:2022, arXiv:KARIMI:2022]. 55.56% of the works use both predictive and ranking quality metrics [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023, IJ:ESWA:MEI:2022, IJ:AI:GAO:2023, IJ:IS:Chaudhuri:2022, IJ:mathematics:GUO:2024, IC:KSEM:YU:2022, IJ:ASC:Wang:2024, IJ:AS:Wang:2024, IJ:ESWA:XIAO:2023, IJ:SMC:WANG:2022, IJ:JOI:Yadav:2022, IJ:Scientometrics:Jiang:2023, ARXIV:Shen:2024, IJ:DSM:Sivasankari:2023, AC:ACIIDS:PAN:2024, IJ:IS:PINEDO:2024, IJ:Scientometrics:ALI:2022, J:EIJ:Huang:2024, IJ:KIS:STERGIOPOULOS:2024, IJ:ESWA:XI:2023, IJ:TIIS:GUO:2024, IJ:SC:XI:2024, IC:EITCE:NIU:2023, IC:MLBDBI:GUAN:2022, IC:ICITACEE:Switrayana:2022, IJ:Scientometrics:Huang:2024, IJ:AI:Thierry:2023, IJ:IEEE:HADHIATMA:2023, IJ:Scientometrics:Thierry:2023, IJ:CDS:AHMEDI:2022, IC:ASSP:Shen:2024, IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023, IJ:Scientometrics:TAN:2023]. 31.75% of the works use only predictive metrics [IJ:DATABASE:Kart:2022, J:JIS:MUNGEN:2022, IJ:KBS:Xiao:2023, IJ:SECS:Hamisu:2024, IJ:JES:LI:2024, IJ:AI:Xiao:2023, IJ:AMEC:KUS:2023, IJ:Heliyon:LI:2024, IJ:JS:Huang:2023, IJ:Scientometrics:GUNDOGAN:2022, PP:Mahesh:2023, IJ:EIT:Smail:2023, IJ:IEEE:STERLING:2022, IJ:ELECTRONICS:Blazevic:2023, IJ:IEEE:REN:2022, ARXIV:Mohamed:2022, arXiv:KARIMI:2022, IC:SEAI:Wang:2024, CF:CSCW:LI:2022, IJ:SHTI:GUO:2022]. Lastly, 11.11% of the works employ only ranking metrics [IJ:Scientometrics:Kanwal:2024, IC:ICDE:XIE:2022, IJ:IPM:Long:2024, IJ:IEEE:SARWAR:2021, IS:CSET:WU:2022, IJ:AS:Chen:2023, EC:ECIR:Takahashi:2022].
@@
[IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023]
[IJ:Scientometrics:TAN:2023]
[IJ:KBS:Xiao:2023]
[IJ:AI:Xiao:2023, IJ:SECS:Hamisu:2024, IJ:JES:LI:2024, IJ:DATABASE:Kart:2022, IJ:AMEC:KUS:2023, J:JIS:MUNGEN:2022]
[IJ:ESWA:MEI:2022]
[IJ:AI:GAO:2023]
[IJ:IS:Chaudhuri:2022, IC:KSEM:YU:2022, IJ:ASC:Wang:2024, IJ:mathematics:GUO:2024]
[IJ:JOI:Yadav:2022]
[IJ:ESWA:XIAO:2023, IJ:SMC:WANG:2022]
[ARXIV:Shen:2024, IJ:Scientometrics:Jiang:2023, IJ:DSM:Sivasankari:2023]
[IJ:EIT:Smail:2023, IJ:Heliyon:LI:2024, IJ:Scientometrics:GUNDOGAN:2022, PP:Mahesh:2023, IJ:JS:Huang:2023]
[IJ:Scientometrics:ALI:2022]
[IJ:KIS:STERGIOPOULOS:2024, IJ:SC:XI:2024, IJ:TIIS:GUO:2024, IJ:ESWA:XI:2023, J:EIJ:Huang:2024, IC:MLBDBI:GUAN:2022]
[IJ:IEEE:HADHIATMA:2023, IJ:Scientometrics:Huang:2024, IJ:Scientometrics:Thierry:2023, IJ:AI:Thierry:2023]
[IJ:CDS:AHMEDI:2022]
[IC:ICITACEE:Switrayana:2022, IC:EITCE:NIU:2023]
[arXiv:KARIMI:2022, ARXIV:Mohamed:2022]
[IC:ASSP:Shen:2024, IJ:IEEE:REN:2022]
[IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023]
[IJ:Scientometrics:Kanwal:2024, IC:ICDE:XIE:2022]
[IJ:AS:Chen:2023, IS:CSET:WU:2022]
[IJ:IEEE:SARWAR:2021]
[IJ:ELECTRONICS:Blazevic:2023, IJ:SHTI:GUO:2022]
[CF:CSCW:LI:2022, IC:SEAI:Wang:2024]
[IJ:AS:Wang:2024]
[IJ:IPM:Long:2024]
[EC:ECIR:Takahashi:2022]
[IJ:IS:PINEDO:2024]
[AC:ACIIDS:PAN:2024]
[IJ:IEEE:STERLING:2022]
@@
In addition to the previously mentioned metrics, other metrics have also been observed in 25.40% of the works, with Area Under the Curve (AUC) and Hit Rate (HR) being the most commonly used. AUC evaluates the ability of the system to distinguish between relevant and irrelevant papers, while HR is the ratio of the number of relevant documents in the Top-N recommendation list to the total number of relevant documents in the dataset. Another, less known metric, bpref, has been used in one work [IJ:Scientometrics:TAN:2023]; it measures retrieval effectiveness by counting how often non-relevant documents are ranked before relevant ones using binary relevance judgments.
@@
AUC [IJ:IEEE:REN:2022, IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023, IJ:KBS:Xiao:2023, CF:CSCW:LI:2022, IJ:SHTI:GUO:2022] 9.52%
HR [IC:ASSP:Shen:2024, IJ:AS:Chen:2023, IJ:AS:Wang:2024, IS:CSET:WU:2022] 6.35%
ACC [IJ:KBS:Xiao:2023, IC:SEAI:Wang:2024] 3.17%
Success [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023] 3.17%
Scalability [IJ:JOI:Yadav:2022] 1.59%
Recency [IJ:JOI:Yadav:2022] 1.59%
High citation popularity [IJ:JOI:Yadav:2022] 1.59%
Diversity [IJ:JOI:Yadav:2022] 1.59%
True Negative Rate (TNR) [IJ:KBS:Xiao:2023] 1.59%
MSE [IJ:KBS:Xiao:2023] 1.59%
bpref [IJ:Scientometrics:TAN:2023] 1.59%
@@ 3.3.3 How are the evaluations conducted?
76.19% of the analyzed papers focus on quantitative assessments. 31.25% of them are complemented by both, ablation studies and configuration tests [IJ:mathematics:GUO:2024, IC:KSEM:YU:2022, IJ:ASC:Wang:2024, IJ:ESWA:XIAO:2023, IJ:SMC:WANG:2022, IJ:Scientometrics:Jiang:2023, ARXIV:Shen:2024, IJ:Scientometrics:ALI:2022, J:EIJ:Huang:2024, IJ:Scientometrics:Huang:2024, IJ:AI:Thierry:2023, IJ:IEEE:HADHIATMA:2023, IC:DASFAA:Wang:2022, IC:ICDE:XIE:2022, IS:CSET:WU:2022]. 29.17% are complemented by ablation studies [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023, IJ:AS:Wang:2024, IJ:JOI:Yadav:2022, IJ:DSM:Sivasankari:2023, IC:EITCE:NIU:2023, IC:MLBDBI:GUAN:2022, IJ:Scientometrics:Thierry:2023, IC:ASSP:Shen:2024, IJ:KBS:Xiao:2023, IJ:JS:Huang:2023, IC:SEAI:Wang:2024, CF:CSCW:LI:2022, IJ:AS:Chen:2023] and 10.42% by configuration tests [AC:ACIIDS:PAN:2024, IJ:TIIS:GUO:2024, IJ:Heliyon:LI:2024, IJ:IEEE:REN:2022, IJ:KIS:STERGIOPOULOS:2024]. Additionally, 8.33% of the works that conduct exclusively quantitative evaluations also assess execution time [J:EIJ:Huang:2024, IC:SEAI:Wang:2024, IJ:KIS:STERGIOPOULOS:2024, IJ:DATABASE:Kart:2022].
14.29% of the works carry out both quantitative and qualitative evaluation, of which 44.44% [IJ:ESWA:MEI:2022, IJ:ESWA:Wang:2023, IJ:SC:XI:2024, IJ:Scientometrics:TAN:2023] complement it with some type of ablation study and 44.44% [IJ:AI:GAO:2023, IJ:IS:Chaudhuri:2022, IJ:ESWA:XI:2023, IJ:IPM:Long:2024] with an ablation study and configuration test.
9.52% of the works conduct only a qualitative evaluation [IC:WIST:Roßner:2023, IJ:IS:PINEDO:2024, J:JIS:MUNGEN:2022, IJ:ELECTRONICS:Blazevic:2023, IJ:Scientometrics:Kanwal:2024, IJ:IEEE:SARWAR:2021], only one includes an additional analysis [IJ:IS:PINEDO:2024] to evaluate execution time and only 2 conduct online evaluations [IJ:IS:Chaudhuri:2022, J:JIS:MUNGEN:2022].
@@
[IJ:AI:GAO:2023, IJ:ESWA:XI:2023, IJ:IPM:Long:2024]
[IJ:ESWA:MEI:2022, IJ:ESWA:Wang:2023, IJ:Scientometrics:TAN:2023]
[IC:ICITACEE:Switrayana:2022]
[ARXIV:Shen:2024, IC:KSEM:YU:2022, IC:DASFAA:Wang:2022, IJ:Scientometrics:Huang:2024, IJ:ESWA:XIAO:2023, IJ:SMC:WANG:2022, IJ:Scientometrics:Jiang:2023, IJ:Scientometrics:ALI:2022, IJ:IEEE:HADHIATMA:2023, IJ:ASC:Wang:2024, IJ:mathematics:GUO:2024, IJ:AI:Thierry:2023, IS:CSET:WU:2022, IC:ICDE:XIE:2022]
[J:EIJ:Huang:2024]
[IJ:AS:Chen:2023, IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023, IJ:JOI:Yadav:2022, IJ:Scientometrics:Thierry:2023, IC:ASSP:Shen:2024, IJ:KBS:Xiao:2023, CF:CSCW:LI:2022, IJ:AS:Wang:2024, IJ:DSM:Sivasankari:2023, IC:MLBDBI:GUAN:2022, IJ:JS:Huang:2023, IC:EITCE:NIU:2023]
[IJ:Heliyon:LI:2024, IJ:IEEE:REN:2022, IJ:TIIS:GUO:2024, AC:ACIIDS:PAN:2024]
[IJ:DATABASE:Kart:2022]
[IJ:AI:Xiao:2023, IJ:Scientometrics:GUNDOGAN:2022, IJ:CDS:AHMEDI:2022, IJ:EIT:Smail:2023, IJ:IEEE:STERLING:2022, PP:Mahesh:2023, arXiv:KARIMI:2022, ARXIV:Mohamed:2022, IJ:SECS:Hamisu:2024, IJ:JES:LI:2024, IJ:SHTI:GUO:2022, IJ:AMEC:KUS:2023, EC:ECIR:Takahashi:2022]
[IJ:Scientometrics:Kanwal:2024]
[IJ:SC:XI:2024]
[IJ:IS:Chaudhuri:2022]
[IC:WIST:Roßner:2023]
[IJ:ELECTRONICS:Blazevic:2023]
[IJ:KIS:STERGIOPOULOS:2024]
[IJ:IEEE:SARWAR:2021]
[J:JIS:MUNGEN:2022]
[IJ:IS:PINEDO:2024]
[IC:SEAI:Wang:2024]
@@
Only 12.70% of the works do not compare their systems or algorithms with others [IJ:Heliyon:LI:2024, IJ:SHTI:GUO:2022, IJ:AI:Xiao:2023, IJ:DATABASE:Kart:2022, IC:WIST:Roßner:2023, IJ:IEEE:SARWAR:2021, IJ:ELECTRONICS:Blazevic:2023, IJ:IS:PINEDO:2024], 50% of which rely solely on qualitative evaluations.
@@
Regarding the types of baselines employed, 28.57% of the works utilized a combination of RP-RS and RSs baselines [IC:MLBDBI:GUAN:2022, IS:CSET:WU:2022, IJ:AS:Chen:2023, IJ:KBS:Xiao:2023, IJ:Scientometrics:Jiang:2023, IJ:ASC:Wang:2024, IJ:mathematics:GUO:2024, J:EIJ:Huang:2024, IC:IKT:JAFARI:2021, IC:ICDE:XIE:2022, IJ:IS:Chaudhuri:2022, IJ:AUT:Jafari:2023, IJ:ESWA:XI:2023, IJ:ESWA:XIAO:2023, IJ:JS:Huang:2023, PP:Mahesh:2023, IJ:SC:XI:2024, IJ:TIIS:GUO:2024], followed by 25.40% that employed only RP-RS baselines [IJ:ESWA:MEI:2022, IJ:JOI:Yadav:2022, IJ:AI:GAO:2023, IJ:Scientometrics:Huang:2024, arXiv:KARIMI:2022, IC:ICITACEE:Switrayana:2022, IJ:CDS:AHMEDI:2022, IJ:IEEE:REN:2022, IJ:Scientometrics:ALI:2022, IJ:Scientometrics:GUNDOGAN:2022, IJ:AI:Thierry:2023, IJ:DSM:Sivasankari:2023, IJ:EIT:Smail:2023, IJ:Scientometrics:Thierry:2023, IJ:Scientometrics:Kanwal:2024, IJ:SECS:Hamisu:2024], and 17.46% that used RSs baselines [ARXIV:Mohamed:2022, CF:CSCW:LI:2022, IC:DASFAA:Wang:2022, IC:KSEM:YU:2022, IJ:SMC:WANG:2022, IC:EITCE:NIU:2023, IJ:ESWA:Wang:2023, IC:ASSP:Shen:2024, IJ:AS:Wang:2024, IJ:JES:LI:2024, IJ:KIS:STERGIOPOULOS:2024]. Notably, 15.87% of the works did not compare recommendation algorithms but rather focused on paper and user representation algorithms [AC:ACIIDS:PAN:2024, ARXIV:Shen:2024, EC:ECIR:Takahashi:2022, IC:SEAI:Wang:2024, IJ:AMEC:KUS:2023, IJ:IEEE:HADHIATMA:2023, IJ:IEEE:STERLING:2022, IJ:IPM:Long:2024, J:JIS:MUNGEN:2022, IJ:Scientometrics:TAN:2023]. Although the analyzed time period is relatively brief, the data suggest a potential shift in trend towards the latter.
Regarding datasets, the majority of works (79.37%) use between one and two datasets. More specifically, 47.62% use only one dataset for the evaluation [IJ:Heliyon:LI:2024, IJ:SHTI:GUO:2022, IJ:DATABASE:Kart:2022, IC:WIST:Roßner:2023, IJ:IS:PINEDO:2024, AC:ACIIDS:PAN:2024, IJ:KIS:STERGIOPOULOS:2024, IJ:Scientometrics:GUNDOGAN:2022, IJ:IEEE:STERLING:2022, EC:ECIR:Takahashi:2022, IJ:CDS:AHMEDI:2022, IJ:IEEE:REN:2022, IJ:EIT:Smail:2023, IJ:SECS:Hamisu:2024, IJ:JES:LI:2024, IJ:Scientometrics:Kanwal:2024, ARXIV:Shen:2024, IC:KSEM:YU:2022, IC:IKT:JAFARI:2021, CF:CSCW:LI:2022, IJ:DSM:Sivasankari:2023, PP:Mahesh:2023, IJ:AMEC:KUS:2023, ARXIV:Mohamed:2022, IJ:AUT:Jafari:2023, IC:EITCE:NIU:2023, IJ:Scientometrics:Jiang:2023, IJ:SMC:WANG:2022, IJ:AS:Wang:2024, IJ:mathematics:GUO:2024], while 31.75% use two datasets [IJ:AI:Xiao:2023, IC:ICITACEE:Switrayana:2022, arXiv:KARIMI:2022, IJ:TIIS:GUO:2024, IJ:AI:GAO:2023, IC:DASFAA:Wang:2022, IJ:Scientometrics:Huang:2024, IC:SEAI:Wang:2024, IS:CSET:WU:2022, IJ:AI:Thierry:2023, IJ:Scientometrics:ALI:2022, IJ:Scientometrics:Thierry:2023, IJ:JS:Huang:2023, IC:ASSP:Shen:2024, J:EIJ:Huang:2024, IJ:ESWA:MEI:2022, IJ:IPM:Long:2024, IJ:ASC:Wang:2024, IJ:IS:Chaudhuri:2022]. An additional 11.11% use three datasets [IJ:AS:Chen:2023, IC:MLBDBI:GUAN:2022, IJ:ESWA:XIAO:2023, IJ:JOI:Yadav:2022, IC:ICDE:XIE:2022, IJ:ESWA:Wang:2023, IJ:Scientometrics:TAN:2023]. It is worth noting that works utilizing more than four or eight datasets to evaluate their results typically rely on a base dataset, building different versions of these to assess specific aspects. In [IJ:IEEE:HADHIATMA:2023], up to four versions of each dataset are generated by randomly selecting 25%, 50%, 75%, and 100% of the data, while in [IJ:SC:XI:2024, IJ:KBS:Xiao:2023], two versions of each base dataset are created to obtain dense and sparse versions.
@@ 3.4 What are the main challenges addressed by the RP-RSs?
Previous surveys in the field identified several open issues in their reviews [IJ:DL:BEEL:2015, IJ:IEEE:BAI:2019, IJ:DL:KREUTZ:2022]. In the rapidly evolving field of RP-RSs   significant advances have been made in recent years. However, despite these strides, several challenges persist that hinder the full realization of potential benefits.
@@
The challenge of disregarding user modeling in RP-RSs persists, as many systems rely on generic texts or specific papers provided by users, rather than incorporating personalized user profiles. This approach assumes that the information needs of the user are implicit and not tailored to individual preferences, which limits the overall effectiveness and relevance of recommendations. Personalization, which adapts the recommendations to the unique needs and behaviors of each user, plays a critical role in improving user satisfaction. Compared to [IJ:DL:KREUTZ:2022], the current work shows a 10% increase in the adoption of user modeling, with 61.90% of the reviewed works now addressing this issue and aiming to offer more personalized recommendations.
@@
Relying solely on precision as the primary evaluation metric in RP-RSs overlooks key factors such as recall, ranking, novelty, diversity, and serendipity. While precision measures the accuracy of the top recommendations, recall focuses on completeness, and ranking metrics like NDCG, MAP, and MRR assess how well the system orders relevant papers. However, other important aspects such as novelty (the ability to suggest new and unexpected papers), diversity (offering a broad range of topics), and serendipity (providing surprisingly relevant recommendations) are often under-explored due to their complexity and rarity in use. These factors are crucial for ensuring that recommendations are not only relevant but also engaging and broad in scope, contributing to a richer user experience. Compared to [IJ:DL:KREUTZ:2022], there has been a 25.65% increase in the use of recall and a 18.12% increase in NDCG, but the adoption of less common metrics has decreased by 21.21%, with AUC and HR being the most widely used metrics at 9.52% and 6.35%, respectively.
@@
In the field of RP-RSs serendipity refers to the process of suggesting unexpected yet valuable papers, thereby facilitating the discovery of novel ideas. However, challenges such as balancing novelty with relevance, assessing the impact of serendipitous recommendations, and personalizing the recommendation process have led to serendipity being underemphasized in recent literature, which has predominantly focused on improving accuracy and personalization. Despite these challenges, serendipity holds significant potential in fostering innovation and expanding the scope of research. Among the works reviewed, only one paper explicitly addresses the concept of serendipity [IJ:IEEE:HADHIATMA:2023]. To address this challenge, in [IJ:IEEE:HADHIATMA:2023] the authors propose a citation recommendation framework that integrates multi-topic communities into a modified version of PageRank.
@@
[IJ:DL:KREUTZ:2022] discusses the balance between relevance and diversity in RP-RSs, highlighting that focusing only on relevance can lead to similar recommendations. Diversity in the generated recommendations can better satisfy the needs of the users. The use of MRR as a quality metric reflected the recognition of this issue, as MRR helps ensure that the suggestions are both relevant and diverse. A small number of recent works have specifically targeted the diversity issue in RP-RSs. [AC:ACIIDS:PAN:2024] identifies a new feature capturing the diversity of research interests, moving beyond the recency of publications to model varied researcher interests. [IJ:KIS:STERGIOPOULOS:2024] introduces a graph of FoS, which uncovers latent knowledge paths to enhance recommendation diversity by discovering relationships in scientific areas. [IJ:JOI:Yadav:2022] proposes measuring diversity by calculating the sum of shortest paths between pairs of items in the recommendation list, helping quantify diversity. [IJ:mathematics:GUO:2024] focuses on enhancing diversity by utilizing indirect social ties within academic teams, narrowing down candidate papers based on shared preferences. [IJ:IS:Chaudhuri:2022] promotes diversity through a Knowledge Discovery Model (KDM), which evaluates paper relevance by the degree of keyword association, aiming to provide diverse topics and sub-topics. Lastly, [IJ:AI:GAO:2023] employs a heterogeneous network and RippleNet to explore potential connections and expand user interests, enhancing the diversity of the recommendations. In the current survey, a small increment in the use of the MRR metric has been detected, making it the fourth most used metric, with 30.16% of the works employing it.
@@
Many works in RP-RSs fail to provide sufficient documentation, such as details about experimental setups, algorithms, or data sources. This lack of transparency makes it difficult for other researchers to replicate or fully understand the methods and findings. A clear indication that this issue remains unresolved is that 21 works were excluded from further analysis in this survey due to missing essential information. It is worth noting that while [IJ:DL:KREUTZ:2022] did not identify any works sharing their code, this survey has found 6 works that have made their developed code accessible [IJ:IS:Chaudhuri:2022, IJ:DATABASE:Kart:2022, IJ:IEEE:SARWAR:2021, AC:ACIIDS:PAN:2024, IJ:KIS:STERGIOPOULOS:2024, IJ:Scientometrics:TAN:2023].